### v2

@@
- Score-based Generative Model for Epeech Enhancement
@@
Invited paper for the SPM Special entitled ”Model-based and Data-Driven Audio Signal Processing”.
@@ Introduction
Generative models follow a different learning paradigm, namely estimating and sampling from an unknown data distribution. This can be used to infer a measure of uncertainty for their predictions and to allow the generation of multiple valid estimates instead of a single best estimate as in predictive approaches [5]. Furthermore, incorporating prior knowledge into generative models can guide the learning process and enforce desired properties about the learned distribution. In particular, diffusion models [6, 1] have emerged as a distinct class of deep generative models that boast an impressive ability to learn complex data distributions such as that of natural images [6, 1], music [7], and speech [8]. Diffusion models generate data samples through iterative transformations, transitioning from a Gaussian prior distribution to a target data distribution, as visualized in Figure 1. This iterative generation scheme is formalized as a stochastic process and is parameterized with a DNN that is trained to address a Gaussian denoising task.
@@
This discrete-time Markov chain formulation of diffusion models can be generalized to continuous-time stochastic processes by letting the number of steps grow infinitely, conversely making the distance between steps infinitely small. This facilitates the design of novel diffusion processes and allows the use of more flexible sampling schemes [1]. Specifically, the corresponding forward diffusion process is defined as a stochastic process , i.e. a collection of random variables indexed by a continuous process time [11]. The process time in stochastic processes intuitively corresponds to the index in Markov chains. It is important to note that the process time is completely unrelated to the time dimension of the audio signal. A single random realization of the stochastic process is depicted by the green trajectory in Figure 1. The conditional distribution characterizing the forward diffusion model is the transition kernel instinctively related to the probability in Markov chains. The transition kernel can be computed by solving a stochastic differential equation (SDE), which is a differential equation where some of the coefficients are random [11]. Specifically, we define the so-called forward SDE as
@@
where the function is referred to as the drift coefficient and relates to the deterministic part of the SDE. The function is called the diffusion coefficient and controls the amount of randomness in the SDE. More precisely, the diffusion coefficient scales the noise injected by the stochastic process . In most cases, is chosen to be a Wiener process, which is a stochastic process with independent and normally distributed increments, i.e. [11]. If the SDE coefficients and are affine with respect to , then the transition kernel has a simple Gaussian form
@@
where the mean and standard deviation can be obtained by solving the SDE and computing the first and second moments of the solution [11].
@@
The reverse diffusion process, i.e. the generation process, is also a stochastic process parameterized by the time flowing in the reverse direction, with . Reversing the time axis in (1) results in another SDE called the reverse SDE whose marginal distributions match those of the forward SDE [1]. Therefore, denoising the sample boils down to solving the reverse SDE
@@
where as the process axis is traveled in the reverse direction. The stochastic process is another Wiener process associated to this reverse process axis, i.e. . The quantity (with representing the gradient operator with respect to ) is called score function and is a vector field pointing toward the local maximum of the logarithmic probability density of the stochastic process state . The score function is generally intractable and we need to approximate it with a DNN called score model . Vincent et al. [12] have shown that the score model can be optimized using denoising score matching, i.e. matching the score of the Gaussian transition kernel instead of the score of the unknown probability . The score of the transition kernel can be obtained from Eq. (2) as
@@
where a data example is first sampled from the training set. Then, a process time is sampled uniformly between and , and the diffusion state is obtained by sampling from the transition kernel (2). Here is a time-dependent scaling factor, chosen empirically to stabilize training [1].
@@
Once the score model has been trained, it allows the generation of new samples from the learned data distribution by solving the reverse SDE (3). In practice, this is done by first discretizing the process time axis into steps with a step size , often chosen uniformly. Then an initial condition is sampled and the reverse SDE (3) is integrated between and using a numerical approximation method called SDE solver [13]. A differential equation solver approximates the trajectory between successive steps and as a piecewise polynomial function (linear if first-order solver, quadratic if second-order, etc.), whose coefficients depend on the terms in Eq. (3). An SDE solver, in particular, considers two polynomial functions (with potentially distinct degrees) to model the deterministic and stochastic terms, respectively. For instance, the widespread Euler-Maruyama method is an SDE solver with first-order polynomial approximation for both the deterministic and stochastic components [13]. The generation process is summarized in the algorithm in Fig. 2.
@@
Deterministic sampling can also be used in place of stochastic sampling by deactivating the randomness source, i.e. removing the Wiener process in (3), and adapting the diffusion coefficient . This turns the reverse SDE in a so-called probability flow ODE [1]. A comparison between stochastic and deterministic sampling is presented in Figure 2. We display three realizations of the stochastic and deterministic sampler. Note that because of the stochastic noise injected at each step, two stochastic sampler trajectories starting at the same initial state may end up reaching two different modes of the target data distribution, whereas the corresponding mean trajectory systematically reaches the same mode. This suggests that a stochastic sampler could be used to obtain more diverse samples and also to improve mode coverage, i.e. better represent the modes of the data distribution regardless of the initial state.
@@
Diffusion processes can be defined for various data representations, depending on the audio application considered. Early works such as [9, 8, 14] directly use the waveform representation, whereas some speech enhancement approaches employ the complex short-time Fourier transform (STFT) domain [15, 10, 16], and several music restoration works consider the Constant-Q Transform (CQT) domain, which is a natural space for harmonic music signals [17, 7, 18]. Learned domains like, e.g., auto-encoder latent spaces, can also be exploited for diffusion to reduce the dimensionality of the original audio data or leverage auto-encoding properties, which gives birth to latent diffusion models [19]. Figure 3 offers a schematic overview of diffusion models defined in various domains.
@@
It should be noted that a connection between score-based diffusion models and continuous normalizing flows has been drawn by Lipman et al., creating a new category of so-called flow matching models [20]. Flow matching methods generalize Gaussian diffusion models and allow to design more flexible probability paths (based on e.g. optimal transport) between arbitrary terminal distributions. This approach has been used to train a foundational speech model, which can be finetuned to perform restoration tasks such as speech enhancement and separation [21].
@@
In statistical model-based speech enhancement, each time-frequency bin of the speech and noise spectrograms is often assumed to be mutually independent and to follow a zero-mean complex Gaussian prior distribution [3]. For an additive mixture model, this yields a Gaussian likelihood model for the mixture and a Gaussian posterior model for the clean speech estimate using Bayes’ rule. Under this Gaussian assumption, the posterior mean can be derived as the celebrated Wiener filter solution, providing the optimal speech estimate in the minimum mean square error (MMSE) sense. However, distributional and independence assumptions are merely approximations utilized out of convenience for the derivation of closed-form estimators, e.g. the mentioned Wiener filter. With diffusion models, there are no distributional and independence assumptions on the speech and noise signals themselves. Indeed, the very intent of deep generative modeling is to allow more flexibility by inferring the signal structure from data rather than the parameters of a fixed distribution.
@@
The second powerful property of diffusion models is their natural integration within stochastic optimization and posterior sampling using Bayes’ theorem, making them particularly suited for conditional generation. We consider the case of audio restoration under the scope of inverse problem solving, i.e. retrieval of clean audio from a measurement . There, an approximation of the measurement likelihood can be obtained via a closed-form model of the operation corrupting into . Combining this likelihood model and the learned deep generative prior with Bayes’ rule can provide sampling or stochastic optimization algorithms for the conditional generation of samples from the posterior distribution [25, 26, 7, 18].
@@
In many restoration tasks such as denoising, dereverberation and separation, the corrupted signal and the clean signal have same dimensionality. This allows to define what we denote as task-adapted diffusion processes, i.e. diffusion processes whose mean is at and at , and that interpolates between these terminal values for . This is a form of conditioning which is not introduced as an auxiliary variable to the score model as for input conditioning, but rather directly injected in the parameters of the diffusion process itself. Examples of classical and task-adapted diffusion processes are visualized on Figure 4. CDiffuSE [9] is one of the earliest methods in this class, formulating the processes in discretized time steps. Score-based Generative Model for Epeech Enhancement (SGMSE) [15] and SGMSE+ [10] extend this idea to the continuous SDE-based formalism of diffusion models to derive pairs of forward and backward processes. Subsequent works [16, 23] build upon this formalism to design alternative forward and backward processes which result in fewer sampling steps and/or higher reconstruction quality. In practice, these methods combine task-adapted diffusion processes with input conditioning, by also providing as an auxiliary input to the score model.
The interpolation between and underlying these approaches assumes an additive signal model typical for denoising tasks, which can also be treated as natural for separation tasks [28] or for convolutive corruptions such as reverberation. The aforementioned methods also achieve excellent reconstruction quality for non-additive corruptions like in bandwidth extension [16] and STFT phase retrieval [29], which shows their ability to perform blind restoration, i.e. when the corruption operator is not perfectly known during inference.
@@
Some works alleviate this issue by simply bypassing the likelihood term, and instead project the state onto the set of the observations at every step of the discretized inference process. [1]. The objective of such projection-based method is to inject the reliable parts of the observations into the intermediate predictions. This ensures that at each step, the intermediate output of the algorithm is consistent with the algorithm input, i.e. the degraded audio, which is often refered to as data consistency and helps avoiding degenerate solutions. Projection-based methods offer the advantage of ensuring data consistency and simplicity in terms of algorithmic implementation. However, their applicability is limited to a reduced set of linear inverse problems, such as audio inpainting or bandwidth extension [19, 7], where closed-form expressions for the projection step are available.
@@
The -norm in (11) can be replaced by any other objective function that better accounts for the statistics of the measurements [25]. For example in [30], the measurement noise is modelled as a Gaussian in the compressed STFT domain, to better account for the heavy-tailedness of speech distributions. It is important to note that the gradient operator requires differentiating through the score model by backpropagation, which introduces a computational overhead. In practice, the unknown measurement noise variance is estimated empirically using e.g. the norm of the gradients in (11) [7]. Compared to projection-based methods, this approach is not limited to linear problems and can be applied to cases where is nonlinear, as long as the operator is differentiable. A geometrical perspective on the sampling process is displayed in Figure 6. This diagram illustrates the intuition behind conditional sampling with a diffusion model, in this case in the context of bandwidth extension. This strategy has been successfully applied in audio bandwidth extension [7], audio inpainting [17], and dereverberation [26].
@@
Reducing per-step inference time: A natural way to accelerate inference is to reduce the cost of each call to the score model. This can be obtained by minimizing the size of the neural network used for score inference through e.g. knowledge distillation, or by reducing the size of the space itself where diffusion is performed, resulting in latent diffusion models. The latent space should be designed such that its reduced dimensionality has a limited impact on the reconstruction quality, and its structure allows for score estimation with a reasonably-sized neural network. Latent diffusion is popular in text-to-audio generation and has been recently applied to audio editing (including restoration) in AUDIT [31], which uses latents provided by a variational auto-encoder.
Improving initialization: Another possibility to accelerate sampling is to find a better initial prediction to reduce the distance between the initial condition and the target sample . This can be provided by a separate plug-in predictive network providing an estimate of the posterior mean as proposed by Lemercier et al. in their Stochastic Regeneration Model (StoRM) [16] for speech enhancement (see Figure 8). The diffusion-based generative model can restore target cues potentially destroyed by the predictive stage while additionally removing residual corruption. The resulting approach requires significantly fewer function evaluations than the original diffusion-only model in [10], for a better-sounding result. Figure 9 shows the clean, degraded, and restored speech spectrograms produced with StoRM. As a simpler alternative, the corrupted utterance can be directly used as the mean of the initial state . This latter strategy is sometimes referred to as warm initialization and has already been used in audio-related tasks such as speech enhancement [9] and bandwidth extension [18]. A good initial prediction can also be obtained by designing a more suitable diffusion trajectory to reduce the mismatch between training and inference, as suggested by Lay et al. [23] for speech enhancement. As shown in Figure 4, the Brownian bridge with exponential diffusion (BBED) SDE proposed in [23] has a linear, constant speed mean interpolating between the clean and noisy speech, which effectively terminates at the clean speech in finite time, unlike the original Ornstein-Uhlenbeck variance-exploding (OUVE) SDE proposed in [15].
@@
Causal processing: In real-time acoustic communications (e.g. hearing aids), future information can not be used to process the current signal which means processing must be causal. Diffusion models can be adapted for causal processing, as in Richter et al. [36], where the convolutional score network architecture and the audio level normalization procedure are modified to meet causality requirements.
@@
Artifacts produced by diffusion models can differ in nature from those produced by statistical signal processing methods or predictive deep learning models. It was observed in [10] that speech enhancement diffusion models tend to hallucinate for negative input signal-to-noise ratios, i.e. when noise dominates clean speech. This can lead to speech inpainting in noise-only regions, breathing and gasping artifacts, or the introduction of phonetic confusion, which may have a negative impact in real-world applications. This behavior can be mitigated by introducing external modalities such as video in Richter et al. [37], where lip movements are analyzed to determine the phoneme used as conditioning for score estimation guidance. Alternatively, as presented in StoRM [16] the input signal-to-noise ratio can be first increased by using a predictive deep learning model to remove parts of the noise, at the potential cost of speech distortions. A generative diffusion model is then used to reconstruct the noisy and distorted speech, which was shown to help avoid hallucination effects and thus increase the robustness to challenging conditions.
Generative pre-training is another approach to increase robustness to outliers. It involves using a pretext task such as masked modeling to train the diffusion model in a self-supervised fashion. Masked modeling involves randomly masking some regions of audio and instructing the model to fill in those masked sections using the available context information, i.e. the non-masked regions. This pre-trained model can then be fine-tuned for a particular downstream task (e.g. speech enhancement, music restoration, etc.) using a supervised setting. Liu et al. [21] show that their diffusion model SpeechFlow benefits from generative pre-training, as it increases its robustness to adverse scenarios such as noise-dominated utterances in speech enhancement. They also notice that generative pre-training consistently increases performance for most speech restoration tasks.

### v3

@@
- Score-based generative models for speech enhancement
@@
Invited paper for the SPM Special entitled “Model-based and Data-Driven Audio Signal Processing”.
@@ Introduction
[…]
@@
This discrete-time Markov chain formulation of diffusion models can be generalized to continuous-time stochastic processes by letting the number of steps grow infinitely, conversely making the distance between steps infinitely small. This facilitates the design of novel diffusion processes and allows the use of more flexible sampling schemes [8]. Specifically, the corresponding forward diffusion process is defined as a stochastic process , i.e., a collection of random variables indexed by a continuous process time [13]. The process time in stochastic processes intuitively corresponds to the index in Markov chains. It is important to note that the process time is completely unrelated to the time dimension of the audio signal. A single random realization of the stochastic process is depicted by the green trajectory traversing Figure 1. The conditional distribution characterizing the forward diffusion model is the transition kernel instinctively related to the probability in Markov chains. The transition kernel can be computed by solving a stochastic differential equation (SDE), which is a differential equation where some of the coefficients are random [13]. Specifically, we define the so-called forward SDE as
@@
where the function is referred to as the drift coefficient and relates to the deterministic part of the SDE. The function is called the diffusion coefficient and controls the amount of randomness in the SDE. More precisely, the diffusion coefficient scales the noise injected by the stochastic process . In most cases, is chosen to be a Wiener process, which is a stochastic process with independent and normally distributed increments, i.e., [13]. If the drift coefficient is an affine function of and the diffusion coefficient is independent of , then the transition kernel has a simple Gaussian form
@@
where the mean and standard deviation are obtained by analytically solving the SDE and computing the first and second moments of the solution [13].
@@
The reverse diffusion process, i.e. the generation process, is also a stochastic process parameterized by the process time flowing in the reverse direction, with . Reversing the process time axis in (1) results in another SDE called the reverse SDE whose marginal distributions match those of the forward SDE [8]. Therefore, denoising the sample boils down to solving the reverse SDE
@@
where as the process axis is traveled in the reverse direction. The stochastic process is another Wiener process associated to this reverse process axis, i.e., . The quantity (with representing the gradient operator with respect to ) is called score function and is a vector field informative about the variations of the process state’s logarithmic probability density. The score function is generally intractable and we need to approximate it with a DNN called score model with parameters . Vincent et al. [14] have shown that the score model can be optimized using denoising score matching, i.e., matching the score of the Gaussian transition kernel instead of the score of the unknown probability . The score of the transition kernel can be obtained from (2) as
@@
where a data point is first sampled from the training set. Then, a process time is sampled uniformly between and , and the diffusion state is obtained by sampling from the transition kernel (2). Here is a time-dependent scaling factor, chosen empirically to stabilize training [8].
@@
Once the score model has been trained, it allows the generation of new samples from the learned data distribution by solving the reverse SDE (3). In practice, this is done by first discretizing the process time axis into steps with a step size , often chosen uniformly. Then an initial condition is sampled and the reverse SDE (3) is integrated between and using a numerical approximation method called SDE solver [15]. A differential equation solver approximates the trajectory between successive steps and as a piecewise polynomial function (linear if first-order solver, quadratic if second-order, etc.), whose coefficients depend on the terms in (3). An SDE solver, in particular, considers two polynomial functions (with potentially distinct degrees) to model the deterministic and stochastic terms, respectively. For instance, the widespread Euler-Maruyama method is an SDE solver with first-order polynomial approximation for both the deterministic and stochastic components [15]. The generation process is summarized in the algorithm in Figure 2.
@@
Deterministic sampling can also be used in place of stochastic sampling by deactivating the randomness source, i.e., removing the Wiener process in (3) and scaling the diffusion coefficient . This turns the reverse SDE into a so-called probability flow ordinary differential equation (ODE) [8]. A comparison between stochastic and deterministic sampling is presented in Figure 2. We display three realizations of the stochastic and deterministic sampler. Note that because of the stochastic noise injected at each step, two stochastic sampler trajectories starting at the same initial state may end up reaching two different modes of the target data distribution, whereas the corresponding mean trajectory systematically reaches the same mode. This suggests that a stochastic sampler could be used to obtain more diverse samples and also to improve mode coverage, i.e., better represent the modes of the data distribution regardless of the initial state.
@@
Diffusion processes can be defined for various data representations, depending on the audio application considered. Early works such as [11, 10, 16] directly use the waveform representation, whereas some speech enhancement approaches employ the complex short-time Fourier transform (STFT) domain [17, 12, 18], and several music restoration works consider the Constant-Q Transform (CQT) domain, which is a natural space for harmonic music signals [19, 9, 20]. Learned domains like, e.g., autoencoder latent spaces, can also be exploited for diffusion to reduce the dimensionality of the original audio data or leverage auto-encoding properties, which gives birth to latent diffusion models [21]. Figure 3 offers a schematic overview of diffusion models defined in various domains.
@@
It should be noted that a connection between score-based diffusion models and continuous normalizing flows has been drawn by Lipman et al., creating a new category of so-called flow matching models [22]. Flow matching methods generalize Gaussian diffusion models and allow to design more flexible probability paths (based on, e.g., optimal transport) between arbitrary terminal distributions. This approach has been used to train a foundational speech model, which can be finetuned to perform restoration tasks such as speech enhancement and separation [23].
@@
In statistical model-based speech enhancement, each time-frequency bin of the speech and noise spectrograms is often assumed to be mutually independent and to follow a zero-mean complex Gaussian prior distribution [2]. For an additive mixture model, this yields a Gaussian likelihood model for the mixture and a Gaussian posterior model for the clean speech estimate using Bayes’ rule. Under this Gaussian assumption, the posterior mean can be derived as the celebrated Wiener filter solution, providing the optimal speech estimate in the minimum mean square error (MMSE) sense. However, distributional and independence assumptions are merely approximations utilized out of convenience for the derivation of closed-form estimators, e.g., the mentioned Wiener filter. With diffusion models, there are no distributional and independence assumptions on the speech and noise signals themselves. Indeed, the very intent of deep generative modeling is to allow more flexibility by inferring the signal structure from data rather than the parameters of a fixed distribution.
@@
The second powerful property of diffusion models is their natural integration within stochastic optimization and posterior sampling using Bayes’ theorem, making them particularly suited for conditional generation. We consider the case of audio restoration under the scope of inverse problem solving, i.e., retrieval of clean audio from a measurement . There, an approximation of the measurement likelihood can be obtained via a closed-form model of the operation corrupting into . Combining this likelihood model and the learned deep generative prior with Bayes’ rule can provide sampling or stochastic optimization algorithms for the conditional generation of samples from the posterior distribution [27, 28, 9, 20].
@@
In many restoration tasks such as denoising, dereverberation and separation, the corrupted signal and the clean signal have same dimensionality. This allows one to define what we denote as task-adapted diffusion processes, i.e., diffusion processes whose mean is at and at , and that interpolates between these terminal values for . This is a form of conditioning which is not introduced as an auxiliary variable to the score model as for input conditioning, but rather directly injected in the parameters of the diffusion process itself. Examples of classical and task-adapted diffusion processes are visualized on Figure 4. CDiffuSE [11] is one of the earliest methods using task-adapted diffusion, formulating the process in discretized time steps. Score-based generative models for speech enhancement (SGMSE) [17] and SGMSE+ [12] extend this idea to the continuous SDE-based formalism of diffusion models to derive pairs of forward and backward processes. Subsequent works [18, 25] build upon this formalism to design alternative forward and backward processes which result in fewer sampling steps and/or higher reconstruction quality. In practice, these methods combine task-adapted diffusion processes with input conditioning, by also providing as an auxiliary input to the score model.
The interpolation between and underlying these approaches assumes an additive signal model typical for denoising tasks, which can also be treated as natural for separation tasks [30] or for convolutive corruptions such as reverberation. The aforementioned methods also achieve excellent reconstruction quality for non-additive corruptions like in bandwidth extension [18] and STFT phase retrieval [31], which shows their ability to perform blind restoration, i.e., when the corruption operator is not perfectly known during inference.
@@
Some works alleviate this issue by simply bypassing the likelihood term, and instead project the state onto the set of the observations at every step of the discretized inference process [8]. The objective of such projection-based method is to inject the reliable parts of the observations into the intermediate predictions. This ensures that at each step, the intermediate output of the algorithm is consistent with the algorithm input, i.e., the degraded audio, which is often refered to as data consistency and helps avoiding degenerate solutions. Projection-based methods offer the advantage of ensuring data consistency and simplicity in terms of algorithmic implementation. However, their applicability is limited to a reduced set of linear inverse problems, such as audio inpainting or bandwidth extension [21, 9], where closed-form expressions for the projection step are available.
@@
The -norm in (11) can be replaced by any other objective function that better fits the statistics of the measurements [27]. For example in [32], the measurement noise is modelled as a Gaussian in the compressed STFT domain, where the compression is a square-root power law on the STFT magnitude, leaving the phase unchanged. This helps accounting for the heavy-tailedness of speech distributions [33]. It is important to note that the gradient operator requires differentiating through the score model by backpropagation, which introduces a computational overhead. In practice, the unknown measurement noise variance is estimated empirically using, e.g., the norm of the gradients in (11) [9]. Compared to projection-based methods, this approach is not limited to linear problems and can be applied to cases where is nonlinear, as long as the operator is differentiable. A geometrical perspective on the sampling process is displayed in Figure 6. This diagram illustrates the intuition behind conditional sampling with a diffusion model, in this case in the context of bandwidth extension. This strategy has been successfully applied in audio bandwidth extension [9], audio inpainting [19], and dereverberation [28].
@@
Reducing per-step inference time: A natural way to accelerate inference is to reduce the cost of each call to the score model. This can be obtained by minimizing the size of the neural network used for score inference through, e.g., knowledge distillation, or by reducing the size of the space itself where diffusion is performed, resulting in latent diffusion models. The latent space should be designed such that its reduced dimensionality has a limited impact on the reconstruction quality, and its structure allows for score estimation with a reasonably-sized neural network. Latent diffusion is popular in text-to-audio generation and has been recently applied to audio editing (including restoration) in AUDIT [34], which uses latents provided by a variational autoencoder.
Improving initialization: Another possibility to accelerate sampling is to find a better initial prediction to reduce the distance between the initial condition and the target sample . This can be provided by a separate plug-in predictive network providing an estimate of the posterior mean as proposed by Lemercier et al. in their Stochastic Regeneration Model (StoRM) [18] for speech enhancement (see Figure 8). The diffusion-based generative model can restore target cues potentially destroyed by the predictive stage while additionally removing residual corruption. The resulting approach requires significantly fewer function evaluations than the original diffusion-only model in [12], for a better-sounding result. Figure 9 shows the clean, degraded, and restored speech spectrograms produced with StoRM. As a simpler alternative, the corrupted utterance can be directly used as the mean of the initial state . This latter strategy is sometimes referred to as warm initialization and has already been used in audio-related tasks such as speech enhancement [11] and bandwidth extension [20]. A good initial prediction can also be obtained by designing a more suitable diffusion trajectory to reduce the mismatch between training and inference, as suggested by Lay et al. [25] for speech enhancement. As shown in Figure 4, the BBED diffusion process proposed in [25] has a linear, constant speed mean interpolating between the clean and noisy speech, which effectively terminates at the clean speech in finite time, unlike the original OUVE diffusion process proposed in [17].
@@
Causal processing: In real-time acoustic communications (e.g., hearing aids), future information can not be used to process the current signal which means processing must be causal. Diffusion models can be adapted for causal processing, as in [39], where the convolutional score network architecture and the audio level normalization procedure are modified to meet causality requirements.
@@
Artifacts produced by diffusion models can differ in nature from those produced by statistical signal processing methods or predictive deep learning models. It was observed in [12] that speech enhancement diffusion models tend to hallucinate for negative input signal-to-noise ratios, i.e., when noise dominates clean speech. This can lead to speech inpainting in noise-only regions, breathing and gasping artifacts, or the introduction of phonetic confusion, which may have a negative impact in real-world applications. This behavior can be mitigated by introducing external modalities such as video in [40], where lip movements are analyzed to determine the phoneme used as conditioning for score estimation guidance. Alternatively, as presented in StoRM [18] the input SNR can be first increased by using a predictive deep learning model to remove parts of the noise, at the potential cost of speech distortions. A generative diffusion model is then used to reconstruct the noisy and distorted speech, which was shown to help avoid hallucination effects and thus increase the robustness to challenging conditions.
Generative pre-training is another approach to increase robustness to outliers. It involves using a pretext task such as masked modeling to train the diffusion model in a self-supervised fashion. Masked modeling involves randomly masking some regions of audio and instructing the model to fill in those masked sections using the available context information, i.e., the non-masked regions. This pre-trained model can then be fine-tuned for a particular downstream task (e.g., speech enhancement, music restoration, etc.) using a supervised setting. Liu et al. [23] show that their diffusion model SpeechFlow benefits from generative pre-training, as it increases its robustness to adverse scenarios such as noise-dominated utterances in speech enhancement. They also notice that generative pre-training consistently increases performance for most speech restoration tasks.